OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4

OpenStudy (anonymous):

3(x+1)=-2(x-1)-4 3x+3=-2x+2-4 3x+3=-2x-2 3x+2x=-2-3 5x=-5 x=-1

OpenStudy (texaschic101):

you can also check your work by subbing in the answer you got.. 3(-1 + 1) = -2(-1-1) - 4 3(0) = -2(-2) - 4 0 = -4-4 0 = 0 (correct) x = -1

OpenStudy (anonymous):

3x+3=-2x+2-4=5x=2-7=5x=-5 therefore x=-1

OpenStudy (anonymous):

@au1 ?